# LVLM-MPC Collaboration for Autonomous Driving: A Safety-Aware and Task-Scalable Control Architecture

### v1

@@ I Introduction
Figure 1: Conceptual comparison of hierarchical LVLM-MPC pipelines. (a) A unidirectional LVLM-to-MPC chain either executes tasks or stalls; (b) a bidirectional variant merely returns a binary infeasibility flag; (ours) a feedback loop, powered by MPC Builder, assembles task-specific OCPs and synthesizes safe recovery maneuvers, enabling seamless task–motion handovers.
@@ III Review of MPC Builder
This study integrates our previously proposed MPC Builder [7] with the LVLM-based task planner. First, we provide a brief overview of MPC Builder, which is a method for generating OCPs for a given task.
@@
$\mathcal{O}_{t}$ $=\bigl{(}U,\sum_{i=1}^{M}\mathcal{P}_{i}\bigr{)}=\bigl{(}U,((\mathcal{P}_{1}\oplus\mathcal{P}_{2})\oplus\cdots)\oplus\mathcal{P}_{M}\bigr{)}.$ (8)
Here, $\mathcal{P}_{i}$ is the $i$-th MPC primitive, and $M$ is the number of MPC primitives. The operator $\oplus$ combines two MPC primitives as
@@ IV-A System Overview
The LVLM task planner and MPC Builder can be executed asynchronously since MPC Builder can continuously execute the last task at a high frequency (with a sufficient control rate) on its own, even without an updated task command from the LVLM. In contrast, the LVLM operates with low frequency, and it is difficult to issue task commands within the vehicle’s control cycle.
@@ IV-D OCP Generation, Feasibility Check, and MPC Switching in MPC Builder
Although this test is not a strict condition, it is a practical way to assess feasibility.
The MPC Switcher then generates an iOCP if $\mathcal{O}_{t,\mathrm{target}}$ is infeasible, assisting the transition as described in Section IV-C. The OCP to be solved, $\mathcal{O}_{t,\mathrm{solve}}$, is chosen as
$\mathcal{O}_{t,\mathrm{task}}\ \text{feasible}\;\Rightarrow\;\begin{cases}\mathcal{O}_{t,\mathrm{solve}}=\mathcal{O}_{t,\mathrm{task}},\\ \mathcal{O}_{\mathrm{prev}}=\mathcal{O}_{t,\mathrm{task}},\\ n_{\mathrm{iOCP}}=0,\end{cases}$ (11)
$\mathcal{O}_{t,\mathrm{task}}\ \text{infeasible}\;\Rightarrow\;\begin{cases}\mathcal{O}_{t,\mathrm{solve}}=\mathrm{iOCP}(\mathcal{O}_{\mathrm{prev}},\mathcal{O}_{t,\mathrm{task}}),\\ n_{\mathrm{iOCP}}=n_{\mathrm{iOCP}}+1.\end{cases}$ (12)
We set an upper limit $n_{\mathrm{max}}$ on consecutive iOCP selections to prevent the system from being trapped in an iOCP. If $n_{\mathrm{iOCP}}>n_{\mathrm{max}}$, MPC Builder rejects the task switch, reverts to $\mathcal{O}_{\mathrm{prev}}$, and marks the task command infeasible, as indicated in Fig. 4. Because the constraints of the iOCP and previous OCP are the same, this reversion is guaranteed to be feasible. This decision is also fed back to the LVLM as a rejection.
@@ V-A Experimental Setup
As the MPC Builder, we implemented six MPC primitives to execute the task commands and ensure safety for surrounding vehicles, as shown in Table I. The Primitive Assigner selects a lateral task primitive based on the task command $T\in\mathcal{T}$ received from the LVLM. It selects a longitudinal task primitive between Constant Speed (CS) and Adaptive Cruise Control (ACC) using a simple rule; CS is chosen when the leading vehicle is located far ahead of the ego vehicle, whereas ACC is chosen when the leading vehicle is as close as $2d_{\mathrm{acc}}$ [m] to the ego vehicle. For safety primitives corresponding to surrounding vehicles, all vehicles in the vicinity of the ego vehicle are selected, up to a maximum of six.
@@ VI Conclusion
This paper proposes an AD planning framework that combines LVLM-based task planning with MPC Builder-based motion planning. Our approach leverages the strengths of both components: the LVLM’s ability to generate high-level task plans and the MPC Builder’s capability to produce safe and task-scalable motion plans. Simulation experiments demonstrated that the proposed method significantly improves safety compared with baseline methods. Because MPC Builder offers high scalability with respect to driving tasks and scenarios, the proposed system has the potential to exploit the LVLM’s versatility without loss of generality. As future work, we will validate the framework in more diverse and complex scenarios as well as in real-world experiments.

### v2

@@ I Introduction
Figure 1: Conceptual comparison of hierarchical LVLM-MPC pipelines. (a) A unidirectional LVLM-to-MPC chain either executes tasks or stalls; (b) a bidirectional variant merely returns a binary infeasibility flag; (ours) a feedback loop, powered by MPC Builder, assembles task-specific Optimal Control Problems (OCPs) and synthesizes safe recovery maneuvers, enabling seamless task–motion handovers.
@@ III Review of MPC Builder
This study integrates our previously proposed MPC Builder [7] with the LVLM-based task planner. First, we provide a brief overview of MPC Builder, which is a method for generating Optimal Control Problems (OCPs) for a given task.
@@
$\mathcal{O}_{t}$ $=\bigl{(}U,\sum_{i=1}^{M}\mathcal{P}_{i}\bigr{)}=\bigl{(}U,((\mathcal{P}_{1}\oplus\mathcal{P}_{2})\oplus\cdots)\oplus\mathcal{P}_{M}\bigr{)},$ (8)
where $\mathcal{P}_{i}$ is the $i$-th MPC primitive, and $M$ is the number of MPC primitives. The operator $\oplus$ combines two MPC primitives as
@@ IV-A System Overview
The LVLM task planner and MPC Builder can run separately. This works because the MPC Builder can keep doing its current task, even when it doesn’t get new instructions from the LVLM. In contrast, the LVLM operates with low frequency, and it is difficult to issue task commands within the vehicle’s control cycle.
@@ IV-D OCP Generation, Feasibility Check, and MPC Switching in MPC Builder
Although this feasibility check is based on the previous predictive states and is not a strict condition, it is a practical way to assess feasibility and has not caused any issues in our experiments.
The MPC Switcher then generates an iOCP if $\mathcal{O}_{t,\mathrm{target}}$ is infeasible, assisting the transition as described in Section IV-C. The OCP to be solved, denoted as $O_{t,\mathrm{solve}}$, is output by the MPC Switcher, which functions according to Algorithm 2. We set an upper limit $n_{\mathrm{max}}$ on consecutive iOCP selections to prevent the system from being trapped in an iOCP. If $n_{\mathrm{iOCP}}>n_{\mathrm{max}}$, MPC Builder rejects the task switch, reverts to $\mathcal{O}_{\mathrm{prev}}$, and marks the task command infeasible, as indicated in Fig. 4. Because the constraints of the iOCP and previous OCP are the same, this reversion is guaranteed to be feasible. This decision is indicated as $\mathrm{is\_rejected}$ in Algorithm 2, and is also fed back to the LVLM as a rejection.
Algorithm 2 MPC Switcher
1:The OCP from the previous time step, $\mathcal{O}_{t,\mathrm{prev}}$, the consecutive execution count of iOCP, $n_{\mathrm{iOCP}}$, and the upper limit of $n_{\mathrm{iOCP}}$, $n_{\mathrm{max}}$
2:function mpc_switcher($\mathcal{O}_{t,\mathrm{task}}$)
3:    if $\mathcal{O}_{t,\mathrm{task}}$ is feasible :
4:         $\mathcal{O}_{t,\mathrm{solve}}\leftarrow\mathcal{O}_{t,\mathrm{task}}$
5:         $\mathcal{O}_{t,\mathrm{prev}}\leftarrow\mathcal{O}_{t,\mathrm{task}}$
6:         $n_{\mathrm{iOCP}}\leftarrow 0$
7:         $\mathrm{is\_rejected}\leftarrow\mathrm{False}$
8:    else
9:         if $n_{\mathrm{iOCP}}<n_{\mathrm{max}}$ :
10:             $\mathcal{O}_{t,\mathrm{solve}}\leftarrow\mathrm{iOCP}(\mathcal{O}_{t,\mathrm{prev}},\mathcal{O}_{t,\mathrm{task}})$
11:             $n_{\mathrm{iOCP}}\leftarrow n_{\mathrm{iOCP}}+1$
12:             $\mathrm{is\_rejected}\leftarrow\mathrm{False}$
13:         else
14:             $\mathcal{O}_{t,\mathrm{solve}}\leftarrow\mathcal{O}_{t,\mathrm{prev}}$
15:             $n_{\mathrm{iOCP}}\leftarrow 0$
16:             $\mathrm{is\_rejected}\leftarrow\mathrm{True}$
17:return $\mathcal{O}_{t,\mathrm{solve}},\mathrm{is\_rejected}$
@@ V-A Experimental Setup
As the MPC Builder, we implemented six MPC primitives to execute the task commands and ensure safety for surrounding vehicles, as shown in Table I. While our implementation focuses on the highway-driving scenario, the MPC Builder framework has been validated across multiple scenarios in [7], showing its potential applicability to diverse tasks and environments.
The Primitive Assigner selects a lateral task primitive based on the task command $T\in\mathcal{T}$ received from the LVLM. It selects a longitudinal task primitive between Constant Speed (CS) and Adaptive Cruise Control (ACC) using a simple rule; CS is chosen when the leading vehicle is located far ahead of the ego vehicle, whereas ACC is chosen when the leading vehicle is as close as $2d_{\mathrm{acc}}\,[\mathrm{m}]$ to the ego vehicle. For safety primitives corresponding to surrounding vehicles, all vehicles in the vicinity of the ego vehicle are selected, up to a maximum of six.
@@ VI Conclusion
This paper proposes an AD planning framework that combines LVLM-based task planning with MPC Builder-based motion planning. Our approach leverages the strengths of both components: the LVLM’s ability to generate high-level task plans and the MPC Builder’s capability to produce safe and task-scalable motion plans. Simulation experiments demonstrated that the proposed method significantly improves safety compared with baseline methods. Because MPC Builder offers high scalability with respect to driving tasks and scenarios, the proposed system has the potential to exploit the LVLM’s versatility without loss of generality.
Our proposed method has several limitations. A primary limitation is the latency of the LVLM. This not only poses a challenge to real-time processing but may also impede immediate responses to abrupt environmental changes. Another limitation is the lack of a systematic method for tuning the cost weights and penalty factors in the MPC. Although this is an inherent challenge in OCPs themselves, the proper tuning of these parameters becomes particularly difficult in complex tasks and scenarios. As future work, we will validate the framework under diverse and complex scenarios and conduct real-world, real-time experiments to clarify and resolve the challenges posed by these limitations.